If the measure of the side of a square is decreased by 3cm, its perimeter is 48cm. What is the measure of its side?
Answer by ankor@dixie-net.com(22740)

If the measure of the side of a square is decreased by 3cm, its perimeter is 48cm.
What is the measure of its side?
:
Let x = the side of the square
then
(x-3) = the decreased side
and
4(x-3) = 48; the decreased square perimeter
divide both sides by 4
x - 3 = 48/4
x - 3 = 12
x = 12 + 3
x = 15 is the measure of it's side